# Awes, Laws, and Flaws From Today’s LLM Research

### v2

@@ Appendix C Extended Results
Table 7: Accuracy for the model for a 95% interval with sample size ($n$). To obtain this interval we sampled i.i.d. about 100 papers and manually labelled them. Given that certain papers did not contain the criteria we evaluated (e.g., the LLM-as-an-evaluator metric is rare in papers prior to 2023), or were too skewed (dialects stands out on this), we sampled extra for these–hence the overcounting in automatic evaluations. Overall, the model performs well as a labeller, although certain criteria were certainly better-performing than others (e.g. open-sourcing versus SOTA claims). We partially attribute this to prompting; but close inspection of the papers noted that the model tended to overlook content, sometimes verbatim, matching the criteria.
Criterion $D_{m,n}$ $D$ $H_{0}$
Statistical Tests 12 10 Accept
Version Declaration 7 8 Reject
Parameter Declaration 8 7 Accept
Account for Randomness 18 16 Accept
Non-English Evaluation 17 11 Accept
Dialect Evaluation 61 50 Accept
Open Sourcing 8 19 Reject
LLM Evaluators 16 26 Reject
Human Evaluators 11 5 Accept
Automatic Evaluators 6 9 Reject
Limitations Sections 8 10 Reject
Ethics Sections 11 12 Reject
Negative Results 17 10 Accept
Error Breakdowns 14 8 Accept
Emergence Claims 10 8 Accept
Can Reason Claims 12 15 Reject
Cannot Reason Claims 25 18 Accept
Super-Human Capability Claims 27 25 Accept
Table 8: Impact of the presence of a given criterion on its citation number. To determine this we split the distribution into samples containing and not containing the criterion. Then we used a Kolmogorov-Smirnov test: the null hypothesis $H_{0}$ decides whether the samples belong to the same distribution. Rejecting $H_{0}$ (when $D_{m,n}<D$) implies that the presence of the criterion does not impact the citation number. Above, rejections occurred for claims of reasoning, automatic and LLM evaluators, declaring the version of the model/API used, as well as the presence of ethics and limitations sections.
Criterion Corpus (%) Subset (%) Stat Sig. Error Breakdown Random. Can Reason Cannot Reason Emergence
Human w/o Automatic 4 16 31 18 12 27 12 35
Human w/o LLM 23 84 35 31 11 29 8 38
Human w/o LLM, Automatic 2 9 29 25 7 36 14 43
LLM w/o Automatic 2 18 35 8 19 19 8 23
LLM w/o Human 8 66 22 23 15 35 4 39
LLM w/o Human, Automatic 0 2 33 0 33 33 0 0
Automatic w/o Human 63 74 22 13 11 27 5 38
Automatic w/o LLM 75 88 26 17 11 27 5 37
Automatic w/o Human, LLM 55 64 22 12 11 26 5 38
Mean (SOTA) 100 100 25 16 10 27 6 39
Table 9: Percentage breakdown of Research Features (experimental protocols) for various subsets of evaluators, for papers claiming SOTA. Highlighted are the metrics that are lower than in the general corpus by $<$ 2%.
@@
Overall, papers claiming SOTA and relying solely in LLM evaluators were exceedingly rare. A considerable amount of papers relied on either only automatic evaluations (55%) or human evaluations without LLMs (23%). Relying on one type of evaluator was rare for humans (2%) and LLMs (statistically insignificant). Of note, a large portion of the papers relying on LLM evaluators without humans (8%) presented error breakdown analyses (23%; compare with automatic evaluation subsets at 13-17%). Claims of reasoning were often done with LLM evaluators and not human evaluators (35%); contrasting with claims that they cannot reason predominantly done with human evaluation alone (14%).
When comparing with the entire corpus (Table 10) we note that SOTA papers use fewer measures of statistical significance across the board. It is also predominant the use of automated metrics (+10% on average) and fewer Human-only and Human and Automatic only metrics (from 6% to 2% and 8% to 4%). On the other hand, LLM evaluators remained steady, showing that most of the papers claiming SOTA used this metric in some form.
Criterion Corpus (%) Subset (%) Stat Sig. Error Breakdown Random. Can Reason Cannot Reason Emergence
Human w/o Automatic 8 28 42 16 14 21 9 35
Human w/o LLM 24 86 38 27 13 22 9 38
Human w/o LLM, Automatic 6 20 38 17 15 20 11 38
LLM w/o Automatic 2 21 49 11 17 26 4 26
LLM w/o Human 7 66 27 21 11 30 5 41
LLM w/o Human, Automatic 0 2 40 0 40 20 0 0
Automatic w/o Human 54 72 29 14 12 23 7 40
Automatic w/o LLM 65 88 31 18 12 22 8 39
Automatic w/o Human, LLM 46 63 29 13 12 22 7 39
Mean 100 100 28 15 11 22 8 41
Table 10: Percentage breakdown of Research Features (experimental protocols) for various subsets of evaluators for the entire corpus. This table is meant to serve as a comparison for the subset of papers claiming SOTA (Table 9).
Criterion 2021-2022 (%) 2022-2023 (%) 2023-2024 (%) $\hat{\beta}$ $R^{2}$
Statistical Tests 12 2 -5 3 0.34
Declaration of Versions -3 9 0 3 0.63
Declaration of Parameters 5 3 -2 2 0.59
Account for Randomness 6 7 -4 3 0.59
Non-English Evaluation -3 2 -3 -1 0.57
Dialect Evaluation -2 1 -1 0 0.51
Open Sourcing -5 -8 -9 -8 0.99
LLM Evaluator 3 5 15 7 0.87
Human Evaluator 0 0 1 0 0.74
Automatic Evaluator 5 -12 10 0 0.0
Limitations Section 22 1 -1 7 0.6
Ethics Section 11 -10 -3 -2 0.14
Negative Results 4 0 -2 1 0.17
Error Breakdowns 3 5 0 3 0.89
Can Reason 9 19 -4 9 0.8
Cannot Reason 3 -1 -2 0 0.0
Emergence 10 4 -7 3 0.31
Super-Human Intelligence Claims 5 1 -2 1 0.52
@@
Table 11: Yearly percentual change (gap) in volume of papers claiming SOTA and presenting the given criterion as an absolute percent. An absolute percent is more interpretable in this scenario: for example, the volume of papers claiming reasoning capabilities in 2023 was 103 out of 294 papers, or 35%. In 2024 this number was 165 out of 535, or 31%. The gap is then -4%. Unlike in Table 12, the percentages for 2024 are not dependent on their recency: 2024 amounts for 46% of the papers evaluated. Overall we observed decreases in all trends, except versioning and the use of all evaluators. We also report the estimated growth rate $\hat{\beta}$ from an ordinary least squares regression (OLS), which may be interpreted as the average change over years; and the coefficient of determination $R^{2}$. When this value is close to $1$, the OLS model is highly confident of the modelled values.
Criterion 2021-22 (%) 2022-23 (%) 2023-24 (%)
Statistical Tests 24 41 -44
Declaration of Versions 59 38 -29
Declaration of Parameters 52 32 -27
Account for Randomness 16 4 17
Evaluation of Non-English Languages 17 25 -40
Dialect Evaluations -1 10 -10
Open Sourcing 102 -35 28
LLMs Evaluators 2 12 29
Humans Evaluators 29 16 13
Automatic Evaluators 2 -37 31
Limitations Sections 56 38 -12
Ethics Sections 56 10 -50
Negative Results 10 31 -29
Error Breakdowns 19 10 -4
Emergence Claims 96 4 -40
Can Reason Claims -74 74 -41
Cannot Reason Claims -3 7 -5
Super-Human Intelligence Claims -7 10 -3
@@
Table 12: Yearly absolute percentage changes for citation ratios (gap) for all our criteria across the years. Highlighted are negative changes. We selected the subset of papers claiming SOTA results, and highlight the papers with gap drops larger than 5%. For example, papers with human evaluators had 7,793 citations in 2022, versus 26,653 without. In 2023, this number was 12,592 (with) and 28,243 (without). The gap is then -54% in 2022 and -38% in 2023, and the change in the gap is +16%. Note that 2024 accounts for most of the papers in our corpus (46%) but has the lowest number of citations (7%): hence, most of the gap drops are expected.

### v3

@@ Appendix B Labeller Reliability
Table 4: Accuracy for the model for a 95% interval with sample size ($n$). Given that some papers did not contain the criteria we evaluated (e.g., the LLM-as-an-evaluator metric is rare in papers prior to 2023), or were too skewed (dialects stands out on this), we sampled extra for these–hence the overcounting in automatic evaluations. Overall, the model performs well as a labeller. Some criteria were better-performing than others (e.g. open-sourcing versus SOTA claims). We attribute this to prompting and model capabilities: close inspection of the papers noted that the model tended to overlook content, sometimes verbatim, matching the criteria.
Appendix C Extended Results
C.1 Corpus Composition
In this section we present the full results of corpus composition results: percentage-wise (Figures 6 and 8) and change over time (Figure 8). We also show our results of the topic analysis work. Most topics had a relatively even distribution (Figure 5), but there was a clear focus on applications and improvements, as opposed to safety or social and environmental impact. This is not indicative of a problem, however, as we noted in Section 4.2 that these topics are on the rise. It is likely that this disparity in volume is propped up by cross-disciplinary applications.
Figure 5: Most common topics in LLM literature, distributed by primary subject. There is a relatively even distribution between cross-disciplinary applications; training, inference, and evaluation methods; applied NLP and general NLP. Within cross-disciplinary applications, these are overwhelmingly in favour of software, and medicine and healthcare.
Figure 6: Arguments Made for the papers claiming SOTA results. Some of the arguments made, especially emergent behaviour, usually showed a lower prevalence of structural features considered to be good research.
Figure 7: Structural Features for the papers claiming SOTA results. Overall we found a low prevalence of papers containing ethics disclaimers, error breakdowns, and negative results.
Figure 8: Research quality for relevant papers claiming SOTA. Certain research features, such as non-English and dialect evaluations are very low, but various research protocols, such as open-sourcing, versioning, and paper declarations are more frequently seen.
C.2 Research Features and Types of Evaluator
@@
Results broken down by Research Feature and type of evaluator are in Figure 9. Papers claiming SOTA and relying solely in LLM evaluators were exceedingly rare. A considerable amount of papers relied on either only automatic evaluations (55%) or human evaluations without LLMs (23%). Relying on one type of evaluator was rare for humans (2%) and LLMs (statistically insignificant). Of note, a large portion of the papers relying on LLM evaluators without humans (8%) presented error breakdown analyses (23%; compare with automatic evaluation subsets at 13-17%). Claims of reasoning were often done with LLM evaluators and not human evaluators (35%); contrasting with claims that they cannot reason predominantly done with human evaluation alone (14%). When comparing with the entire corpus, SOTA papers use fewer measures of statistical significance across the board. It is also predominant the use of automated metrics (+10% on average) and fewer Human-only and Human and Automatic only metrics (from 6% to 2% and 8% to 4%). On the other hand, LLM evaluators remained steady, showing that most of the papers claiming SOTA used this metric in some form.
Figure 9: Percentage breakdown of Research Features (experimental protocols) for various subsets of evaluators, for papers claiming SOTA. Marked with an arrow pointing down are metrics lower than the general corpus by $<$ 2%. When compared to the general corpus, SOTA papers used fewer measures of statistical significance, and relied more in LLM evaluators. That said, papers only LLMs as evaluators were exceedingly rare, and mostly related to papers claiming emergent behaviour.
C.3 Yearly Topic Changes
The yearly topic changes can be found in Figure 10. While there is a decrease on general NLP–likely attributed to the shift from LLMs as research subjects to research tools–there are considerable increases in other areas, such as cross-disciplinary and multimodal applications. Safety and security appeared to also be modestly rising.
Figure 10: Yearly change in topic distributions for papers focusing on LLMs. We noticed a yearly increase in the volume of papers that involved multimodality, security and safety, and cross-disciplinary applications, which coincides with a more widespread adoption of this technology.
C.4 Criteria versus Citations Analysis
To determine whether the presence of a criterion impacted the number of citations, we used a Kolmogorov-Smirnov test. This test is suitable for the task because it is non-parametric, and hence more robust to priors, at the expense of needing larger data sizes.
Concretely, the null hypothesis $H_{0}$ in this test is that both samples come from the same underlying distribution. Accepting (rather, being unable to reject) $H_{0}$ means that the distributions are statistically indistinguishable, and we are unable to conclude that likelihood of citation is impacted by the presence of the criterion. Rejecting $H_{0}$ implies that the distributions are distinct, and we may conclude that the criterion does impact the number of citations received.
Given a calibrated $p$-value and samples of lengths $m$ and $n$, the Kolmogorov-Smirnov test condition is given by
$D_{m,n}=\sqrt{-\ln{(p/2)}\cdot(\frac{n+m}{2mn})}.$ (1)
If the test statistic (percentage citation difference) $D$ is $D>D_{m,n}$, we reject $H_{0}$. We calibrated our $p$ value to $p<0.05$. An interpretable version of this calibration is that we expect to be wrong about our conclusions 5% of the time. We capture our results in Table 5. Overall, we deemed that 7 out of the 18 criteria did not impact the number of citation number. Of note these were the use of LLMs as evaluators, open sourced artifacts, and the presence of ethics and limitations sections.
Criterion $H_{0}$
Statistical Tests Accept
Version Declaration Reject
Parameter Declaration Accept
Account for Randomness Accept
Non-English Evaluation Accept
Dialect Evaluation Accept
Open Sourcing Reject
LLM Evaluators Reject
Human Evaluators Accept
Automatic Evaluators Reject
Limitations Sections Reject
Ethics Sections Reject
Negative Results Accept
Error Breakdowns Accept
Emergence Claims Accept
Can Reason Claims Reject
Cannot Reason Claims Accept
Super-Human Capability Claims Accept
Table 5: Impact of the presence of a given criterion on its citation number. To determine this we split the distribution into samples containing and not containing the criterion and ran a Kolmogorov-Smirnov test. Rejecting $H_{0}$ (in blue) implies that the presence of the criterion does not impact the citation number.
C.5 Yearly Changes
We show yearly trends in the gap for the volume of papers claiming SOTA (Figure 11) and for the citation ratios (Figure 12) per criteria. The core findings of this section are that in terms of volume (recall that 2024 comprises almost half of our corpus), there was a decrease in papers having statistical tests, claims of emergence and reasoning, and open sourcing. There were, however, increases in the use of LLMs as evaluators. Some other criteria remained steady, such as dialect evaluations. This metric is not susceptible to recency bias. On the other hand, there is a recency bias in the citation ratios, since, by definition, this is a time-dependent metric, and so drops in 2024 are expected.
@@ C.5 Yearly Changes
Figure 11: Yearly percentual change (gap) in volume of papers claiming SOTA and presenting the given criterion as an absolute percent. This quantity is more interpretable in this scenario: for example, the volume of papers claiming reasoning capabilities in 2023 was 103 out of 294 papers, or 35%. In 2024 this number was 165 out of 535, or 31%. The gap is then -4%. Unlike in Figure 12, the percentages for 2024 are not dependent on their recency: 2024 amounts for 46% of the papers evaluated. We observed decreases in all trends, except versioning and all evaluators.
@@
Figure 12: Yearly absolute percentage changes for citation ratios (gap) for all our criteria across the years, for papers claiming SOTA. Papers with human evaluators had 7,793 citations in 2022, versus 26,653 without. In 2023, this number was 12,592 (with) and 28,243 (without). The gap is then -54% in 2022 and -38% in 2023, and the change in the gap is +16%. Note that 2024 accounts for most of the papers in our corpus (46%) but has the lowest number of citations (7%) due to recency bias: hence, most of the gap drops are expected.
Appendix D Future Trends in Citation Volume
Our work parted from the assumption that most LLM-related works would cite at least one of the GPT-3 or GPT-4 papers. This assumption must be examined closely, and in particular be put in contrast with future trends like the rise of–comparatively–open models such as LlaMA.
For this, we performed a follow-up study where we examined the citation volume a year after the cutoff date for the data used in this paper. Overall, most models have not reached the citation volume of the original GPT-3.5 or GPT-4 works. However, there exist exceptions to this rule: the first LlaMA paper, released at the same time as the GPT-4 technical report, has over 5,000 more citations than GPT-4’s; but about a third of GPT-3’s. In comparison, most papers have up to a third of GPT-4’s (Gemini, with 4,200; Gemini Team 2025). The LlaMA paper is impactful, however: remark that the GPT-3 paper is three years older. Three years, as per our results, symbolises a significant amount of relative time in this field.
Even though it is very likely some articles could progressively stop citing both GPT papers, we expect this trend to be farther in the future. Likewise, it is very possible that some, if not most, of the citations in LlaMA papers include also references to at least one of the GPT works. Hence, our core assumption holds–to a degree, given the omission of the LlaMA paper in our work.
Unfortunately, at the time of writing this, most public APIs, including the Internet Archive’s Wayback Machine and Publish or Perish, were no longer able to query Google Scholar. This in turn hindered our ability to perform a comprehensive follow-up. Instead, we attach screenshots for the citation trends in Figure 13.
Figure 13: Citation volume from (clockwise from the top left) the LlaMA, Qwen Bai et al. (2023), Gemini, GPT-4, GPT-3, and Deepseek DeepSeek-AI (2025) models, as of 27 May 2025. While most models do not reach the volumes of the papers studied in this work, there is a definite increasing trend in some LLM-related works. In particular, the Llama paper has 5,000 more citations than GPT-4’s technical report, in spite of both papers being released the same year.